The standard time for an operation in Work-for-Pay Manufacturing Company is 0.25 hour per unit. There are 8 working hours in a day. The base (day rate) pay for an operator of this operation is $14 per hour. Based on these information and using a 120% incentive pay system, what is the daily pay (in $ per day) for an operator who produced 45 units in 8 working hours?

a. 112.00

b. 157.50

c. 166.60

d. Not enough information given to answer this question
Answer #1

No. of units per hour = 1 / Hours per unit = 1 / 0.25 = 4 units per hour

Total Base manufacturing rate per 8 hours = No. of units per hour * No. of hours = 4 * 8 = 32 units

Base total pay per 8 hours = Base pay rate * No. of hours = $14 * 8 = $112

Base pay per unit = Base total pay per 8 hours / Total Base manufacturing rate per 8 hours = 112 / 32 = $3.5

No. of units actually produced per 8 hours = 45 units

Incremental units produced = 45 - 32 = 13 units

Pay for this additional units = Base pay rate * Incentive pay % * No. of additional units = $3.5 * 120% * 13 = $54.6

Total pay = Base total pay per 8 hours + Pay for this additional units = 112 + 54.6 = $166.60

Answer: c) 166.60

-----------------------------------------------------------------------------------------------------------------------

In case of any doubt, please ask through the comment section before Upvote/downvote.

If you liked the answer, please give an upvote. This will be quite encouraging for me, thank-you!!!